AIDA
GELINA
BRIKEN
nToF
CRIB
ISOLDE
CIRCE
nTOFCapture
DESPEC
DTAS
EDI_PSA
179Ta
CARME
StellarModelling
DCF
K40
EDI_PSA
Draft saved at 00:00:00
Fields marked with
*
are required
Entry time:
Thu Apr 3 19:29:05 2025
Author
*
:
Subject
*
:
Using a 1x1 Hamamatsu photodiode (uncoated) +75V bias supplied via RAL108 and standard photodiode mount. Imon ~0 Note negative bias would be supplied with standard lemo cable + standard mount. Positive bias supplied via modified lemo cable to read *positive* pulses after RAL108 preamp ->CHANGED dynamic range to 0.5 Vpp (from 2 Vpp) R14 Positive signals 2.5 Hz Pulser 1024 window / 600 pre-gate trigger 70 threshold / 2300 baseline 55098 DC offset 208Bi source DAQ stopped acquiring for unknown reasons at 22:30. Enough data to analyse. Swapped Bi source for 3-alpha. Had to move diode around. Pick-up / microphonic noise much worse now. Baseline too unstable to trigger. Giving bias results in 1 uA / 25 V current (vs. ~0 before). No alphas can be seen. Further investigation needed. After plugging/uplugging wire from diode leakage current no longer grows to 1 uA / 25 V when giving bias. Behaviour appears inconsistent. Heat shrink applied to makeshift connector to improve situation. Now running without leakage current at 75 V. Note 208Bi source was found to be leaky and removed. R15 Positive signals 2.5 Hz Pulser 1024 window / 650 pre-gate trigger 70 threshold / 2300 baseline 55098 DC offset 3-alpha source
Encoding
:
HTML
ELCode
plain
Suppress Email notification
Attachment 1:
Drop attachments here...
Draft saved at 00:00:00
ELOG V3.1.4-unknown